OpenStudy (mrs.ambrose614):

CAN ANYONE HELP ME WITH THSI QUESTION @just_one_last_goodbye Question 1 (Multiple Choice Worth 1 points) (04.05) The population f(x), in millions, of State A of a country after x years is represented by the function shown below: f(x) = 4(1.08)x The graph shows the population g(x), in millions, of State B of the country after x years: Graph of function g of x equals 2 multiplied by 1.08 to the power of x Which conclusion is correct about the population of State A and State B? The original population of State A was double of the original population of State B. The original population of State B was double o CAN ANYONE HELP

OpenStudy (sunnnystrong):

\[g(x)=2(1.08)^x\] B ^^ \[f(x)=4(1.08)^x\] A^^ Okay so recall that --> \[A=A _{0}(b)^x\] Where: Ao= initial x=time b=growth/decay factor A=amount @Mrs.ambrose614 ... what do you think?

OpenStudy (mrs.ambrose614):

Yes so the ANWSER is b???

OpenStudy (mrs.ambrose614):

Is the ANWSER b @sunnnystrong

OpenStudy (sunnnystrong):

@Mrs.ambrose614 ... No it would be: The original population of State A was double of the original population of State Because.... Ao is twice that of state B..